OpenStudy (anonymous):

The perimeter of the rectangle is 44 cm. Find the value of x. 7 cm ███ 3x cm

OpenStudy (anonymous):

10

OpenStudy (lgbasallote):

P = 2(L + W) P = 2(7 + 3x) 44 = 14 + 6x 30 = 6x x = 5